# IllusionVQA: A Challenging Optical Illusion Dataset for Vision Language Models

### v2

@@ Abstract
The advent of Vision Language Models (VLM) has allowed researchers to investigate the visual understanding of a neural network using natural language. Beyond object classification and detection, VLMs are capable of visual comprehension and common-sense reasoning. This naturally led to the question: How do VLMs respond when the image itself is inherently unreasonable? To this end, we present IllusionVQA: a diverse dataset of challenging optical illusions and hard-to-interpret scenes to test the capability of VLMs in two distinct multiple-choice VQA tasks - comprehension and soft localization. GPT4V, the best performing VLM, achieves 62.99% accuracy (4-shot) on the comprehension task and 49.7% on the localization task (4-shot and Chain-of-Thought). Human evaluation reveals that humans achieve 91.03% and 100% accuracy in comprehension and localization. We discover that In-Context Learning (ICL) and Chain-of-Thought reasoning substantially degrade the performance of Gemini-Pro on the localization task. Tangentially, we discover a potential weakness in the ICL capabilities of VLMs: they fail to locate optical illusions even when the correct answer is in the context window as a few-shot example. [^1] We will make the code and data available at https://github.com/csebuetnlp/IllusionVQA.
@@ 1 Introduction
Figure 1: Performance of state-of-the-art VLMs and human evaluators on IllusionVQA-Comprehension task.
Figure 2: Comparison of IllusionVQA with prior illusion datasets - GVIL (Zhang et al., 2023) and HallusionBench (Liu et al., 2023a).
@@
Optical illusions are characterized by a visual percept that arguably differs from reality. Illusions come in a wide variety; their categorization is difficult because the underlying cause is often unclear Bach and Poloschek (2006); Gregory (1997a). Figure 3 shows examples that are distinct from each other but all fall under the umbrella of optical illusions. Figure 3(a) shows an impossible object where - “Parts of the object (in this case the elephant’s legs) become the background, and vice versa." Shepard (1990). Figure 3(b) depicts a 3-by-3 grid where some cells appear pale-yellow, but upon closer inspection, it is evident that all cells are perfectly white. Figure 3(c) shows a realistic 3D drawing of a white cube with a cup positioned as resting as if on top of the fictional cube. Although all three examples are arguably optical illusions, they have little in common. Despite the challenges in optical illusion classification, Gregory (1997b) proposed four main classes based on appearance - ambiguities, distortions, paradoxes, and fictions. Based on this framework, Figure 3(a) is a paradox since the object depicted cannot truly exist. Figure 3(b) is a distortion since none of the regions is pale yellow, although it might appear so. Figure 3(c) is a Trompe-l’œil illustration and its classification as an optical illusion is contentious Wade and Hughes (1999), although it might fall under the aforementioned fiction category.
@@
The cause of our perception of optical illusions is a rich field of research in Cognitive Psychology Helmholtz (1948); Gregory (1968); Gentilucci et al. (1996). Motivated by similarities in human visual cognition and artificial neural networks Cichy et al. (2016); Eickenberg et al. (2017), researchers have explored if and how artificial neural networks perceive optical illusions Gomez-Villa et al. (2019, 2020); Sun and Dekel (2021); Lonnqvist et al. (2021); Hirsch and Tal (2020). However, the scope and generalizability of these works were limited by the need for a case-by-case analysis of the activations and gradients of the model in response to individual optical illusions. As a remedy, Zhang et al. (2023) was the first to explore optical illusions through small VLMs (up to 13 billion parameters) and to investigate their internal beliefs about optical illusions through natural language.
@@
Unlike prior work, we curate challenging optical illusions from the Internet that span 12 distinct categories inherited from cognitive psychology studies. We craft detailed question-answer pairs and several incorrect options designed to probe the abilities of VLMs. We frame the problem as a standard visual question-answering task (VQA) and evaluate state-of-the-art VLMs such as GPT4V Achiam et al. (2023) and Gemini-Pro Team et al. (2023) in addition to smaller open-source models. Our detailed categorization and extensive human evaluation allow for a fine-grained comparison of human and VLM cognition of optical illusions along different axes. Our contributions are summarized as follows.
@@
    We introduce IllusionVQA, a dataset designed to rigorously test the ability of VLMs to locate and comprehend challenging optical illusions.
@@
    We comprehensively test a wide range of open-source and closed-source VLMs. Our experiments reveal that GPT4V is the most capable model on illusion comprehension and localization but still lags significantly behind human-level performance as shown in Figure 1.
@@ 2 Related Work
Researchers have shown that convolutional neural networks trained on ordinary images are susceptible to certain optical illusions, similar to humans Gomez-Villa et al. (2019, 2020); Afifi and Brown (2019); Benjamin et al. (2019); Sun and Dekel (2021). However, most previous experiments have focused on specific categories of optical illusions and their methodologies do not generalize to all types.
@@
Prior research has yielded multiple tools for synthetic optical illusion generation. Hirsch and Tal (2020) and Fan and Zeng (2023) automatically produce color-related and grating-related optical illusions, respectively. Gomez-Villa et al. (2022) explores creating illusions using GANs. The primary limitation of synthetic illusion generation is the limited variety in the illusions they can create, especially compared to the wide spectrum of optical illusions discovered in the literature. As such, relying solely on existing synthetic illusion generation algorithms to evaluate VLMs severely limits the scope of the evaluation.
In contrast to prior works, we collect images of optical illusions from the internet to ensure data diversity and test VLMs along multiple axes. We manually create a multiple-choice question-answering (QA) dataset where each question is crafted to have only one unambiguous answer while all other options appear plausible but are incorrect. Beyond comprehension, we test how well VLMs can locate geometrically impossible objects in an image. Our experiments probe VLM’s dependence on language priors Goyal et al. (2017) and feature entanglement Tang et al. (2023) in the context of optical illusions.
@@ 3 The IllusionVQA Dataset
IllusionVQA is a Visual Question Answering (VQA) dataset with two sub-tasks. The first task tests comprehension on 435 instances in 12 optical illusion categories. Each instance consists of an image with an optical illusion, a question, and 3 to 6 options, one of which is the correct answer. We refer to this task as IllusionVQA-Comprehension. The second task tests how well VLMs can differentiate geometrically impossible objects from ordinary objects when two objects are presented side by side. The task consists of 1000 instances following a similar format to the first task. We refer to this task as IllusionVQA-Soft-Localization. This section details image collection and filtering, followed by QA generation and classification for the Comprehension task, and finally procedural QA generation for the Soft-Localization task.
@@ 3.1 Image Collection and Filtering
We scraped more than 3500 images of optical illusions from multiple online repositories. We manually inspected each image to ensure that they were optical illusions. We then filter out easy-to-understand illusions using GPT4V. Specifically, we asked GPT4V to describe the image and inspect its response. We did not include an optical illusion in IllusionVQA if GPT4V was able to detect and describe it accurately. In some cases, GPT4V was able to detect the presence of an optical illusion but could not describe it (e.g. it mistook an Impossible Cube Penrose and Penrose (1958) as a Necker Cube Rosenholtz (2011)). We included such cases in IllusionVQA. This leaves us with 374 images of high-quality optical illusions, all of which pass the internal filter checks of the latest GPT4V and Gemini-Pro APIs. We highlight a few examples of filtered-out images in Appendix I.
@@ 3.2 QA Generation for IllusionVQA-Comprehension
We pose the comprehension task in a standard multiple-choice VQA setting similar to Gurari et al. (2018); Goyal et al. (2017). In some examples, this involves simply realizing that there is an illusion (Figure 3(a)). In other cases, we ask VLMs specific questions about the image to probe their understanding of illusions (Figure 3(c)).
We construct suitable questions that probe a VLM’s understanding of optical illusions beyond simply detecting their presence. Each question in IllusionVQA has one correct answer and multiple wrong options. We write the questions and the correct answers and include definitively incorrect alternate options, rather than under-specified or ambiguous. We use the following methods and heuristics to generate the incorrect choices:
@@
    Since VLMs are prone to ignoring visual input Goyal et al. (2017); Jabri et al. (2016), we include incorrect descriptions generated by VLMs when prompted with only the question and without the image.
@@
    We manually inspect and edit each incorrect option to ensure that they are challenging, given the question and the image.
We create multiple questions for an image when appropriate and ensure that the questions are distinct and not simply rephrased versions of one another. We use the Perspectives API [^2] https://www.perspectiveapi.com/ to filter out toxic content in the questions, choices, and answers. Altogether, we created 439 question-answer pairs from the 374 carefully selected optical illusion images. Through rigorous filtering and question-answer generation techniques, we ensured that every instance in the IllusionVQA dataset is highly challenging yet unambiguous. Our dataset includes 30 images similar to those present in the GVIL Zhang et al. (2023) and HallusionBench Liu et al. (2023a) datasets but differ in the specific question asked. Furthermore, our dataset evaluates VLMs through multiple-choice QA instead of true-false QA in GVIL and HallusionBench.
@@ 3.3 IllusionVQA-Comprehension Classification
We adapt the classification methods proposed by Gregory (1997b) and Bach and Poloschek (2006) to our specific dataset and classify 12 classes as shown in Figure 4. We opt for an appearance-based classification that better fits our goal of studying VLMs in a black-box setting, without access to weights, gradients, and activations. There is no clear consensus on classifying optical illusions, so some overlap between classes was unavoidable. In particular, the Deceptive Design category could be considered a subset of the Real Scene category and the same goes for Edited Scene for the Impossible Objects category. However, the images were visually distinct enough to warrant separate categorization. Moreover, the Deceptive Design, Edited Scene, and Upside Down categories are not generally considered as optical illusions in Cognitive Science. We include these images because they are often considered optical illusions colloquially. We discuss the relevant literature on optical illusion and its classification in Appendices A and C.
@@ 3.4 Procedural QA Generation for IllusionVQA-Soft-Localization
Figure 5: Examples of IllusionVQA-Soft-Localization
We procedurally combine images of “Impossible Objects" (Figure 4) to test if VLMs can locate illusions within a scene. Although VLMs have shown impressive capability in generating bounding boxes Bai et al. (2023); Wang et al. (2023), GPT4 and Gemini-Pro were not designed for this task. We opt for an easier localization task by stitching together two images of geometrically impossible objects horizontally and prompting the VLMs to locate the impossible object (Figure 5). We refer to this task as Soft-Localization - since the VLMs need only respond with “left" or “right" (as opposed to Precise Localization where a model outputs exact bounding box coordinates)
@@ 4 Experimental Setup
Since open-source VLMs do not support multiple image input, we evaluate open-source VLMs in the 0-shot and closed-source VLMs in both 0-shot and 4-shot settings. For IllusionVQA-Comprehsion, we selected one sample from the four most common illusion categories as a few-shot example. As such, the IllusionVQA-Comprehension test set contains 435 instances (370 images). Due to the inherent difficulty in describing optical illusions in natural language, we did not attempt Chain-of-Thought (CoT) evaluation on the comprehension task.
For IllusionVQA-Soft-Localization, we selected one instance of each of the four alternate choices as a few-shot example. We provide a standardized reasoning template for each example in the case of 4-shot+CoT evaluation. We use the default API arguments and generation parameters.
We resize all images to 512 pixels for all models while maintaining the aspect ratio. For the soft-localization task, we apply a grayscale transform since the image color is not related to the geometry of an object. We used the latest versions of GPT4V and Gemini-Pro available at the time of writing (gpt-4-vision-preview - 5 March 2024 and gemini-pro-vision - 8 March 2024). We tested three open-source VLMs, namely InstructBLIP Dai et al. (2024), LLaVA-1.5 Liu et al. (2023b), and CogVLM Wang et al. (2023).
@@ 5 Results
This section covers the performance of the different VLMs tested on the two tasks in IllusionVQA: Comprehension and Soft Localization. We make several surprising discoveries and conduct further experiments to validate our findings.
@@
Edited-Scene 21 42.86 61.90 42.86 66.67 71.43 66.67 80.95 100
Upside-Down 7 42.86 71.43 71.43 57.14 71.43 57.14 71.43 100
Pos.-Neg. Space 7 57.41 42.86 71.43 85.71 57.14 71.43 85.71 100
Circle-Spiral 6 33.33 0.00 16.67 33.33 50 33.33 33.33 66.67
@@ 5.1 IllusionVQA-Comprehension
GPT4V leads other VLMs in both 0-shot and 4-shot settings by a wide margin, even though it falls far short of human performance. We conducted further case analysis into GPT4V’s output. GPT4V leads in ten of the twelve IllusionVQA categories and shows competitive performance in the remaining two. Categories such as Real-Scene, Size, Color, and Deceptive Design test a VLM’s perception of depth, color, and challenging scenes; skills that are particularly relevant when using VLMs as autonomous robots in the real world Zitkovich et al. (2023); Wake et al. (2023). GPT4V maintains substantial leads in all four categories and is fairly resistant to challenging real scenes, with a 10.93% lead over Gemini-Pro.
@@
We tested GPT4V and Gemini-Pro using 4-shot learning and observed marginal improvements in overall accuracy in IllusionVQA-Comprehension. However, the improvement was not consistent between different categories. For example, 4-shot GPT4V accuracy was lower than 0-shot in three categories, namely Real-Scene, Hidden Illusion, and Circle-Spiral Illusions. Gemini-Pro shows similar accuracy drops but in different categories. It is possible that ICL is not a ubiquitous strategy because few-shot examples introduce additional language priors that bias the VLMs towards incorrect answers Goyal et al. (2017); Jabri et al. (2016).
Performance of VLMs on Ordinary Vertically-Flipped Images
The Upside-Down category for images that depict two or more intertwined entities depending on vertical orientation. As shown in Table 1, the smaller open-source VLMs perform competitively in this category while significantly lagging behind closed-source VLMs in most others. To validate this discrepancy, we curated a subset of VQA-v2.0 Goyal et al. (2017) and tested LLaVA-1.5 and Gemini-Pro on both unedited and vertically flipped images. Table 2 shows that Gemini-Pro’s accuracy drops by 15.5% on flipped images, while LLaVA-1.5’s accuracy only drops by 6.5%, confirming our findings in Table 1.
@@
Table 2: Performance of VLMs on a subset of 200 instances from VQA-v2.0 Goyal et al. (2017). The model outputs were manually evaluated.

### v3

@@ Abstract
The advent of Vision Language Models (VLM) has allowed researchers to investigate the visual understanding of a neural network using natural language. Beyond object classification and detection, VLMs are capable of visual comprehension and common-sense reasoning. This naturally led to the question: How do VLMs respond when the image itself is inherently unreasonable? To this end, we present IllusionVQA: a diverse dataset of challenging optical illusions and hard-to-interpret scenes to test the capability of VLMs in two distinct multiple-choice VQA tasks - comprehension and soft localization. GPT4V, the best performing VLM, achieves 62.99% accuracy (4-shot) on the comprehension task and 49.7% on the localization task (4-shot and Chain-of-Thought). Human evaluation reveals that humans achieve 91.03% and 100% accuracy in comprehension and localization. We discover that In-Context Learning (ICL) and Chain-of-Thought reasoning substantially degrade the performance of Gemini-Pro in the localization task. Tangentially, we discover a potential weakness in the ICL capabilities of VLMs: they fail to locate optical illusions even when the correct answer is in the context window as a few-shot example.111  The code and datasets are available at https://github.com/csebuetnlp/IllusionVQA.
@@ 1 Introduction
Figure 1: Left: Comparison of human and VLM performance in IllusionVQA-Comprehension. Right: Comparison of IllusionVQA with prior illusion datasets - GVIL (Zhang et al., 2023) and HallusionBench (Liu et al., 2023a).
@@
Optical illusions are characterized by a visual percept that arguably differs from reality. Illusions come in a wide variety; their categorization is difficult because the underlying cause is often unclear (Bach & Poloschek, 2006; Gregory, 1997a). Figure 2 shows examples that are distinct from each other but all fall under the umbrella of optical illusions. Figure 2(a) shows an impossible object where - “Parts of the object (in this case the elephant’s legs) become the background, and vice versa." (Shepard, 1990). Figure 2(b) depicts a 3-by-3 grid where some cells appear pale-yellow, but upon closer inspection, it is evident that all cells are perfectly white. Figure 2(c) shows a realistic 3D drawing of a white cube with a cup positioned as resting as if on top of the fictional cube. Although all three examples are arguably optical illusions, they have little in common. Despite the challenges in optical illusion classification, Gregory (1997b) proposed four main classes based on appearance - ambiguities, distortions, paradoxes, and fictions. Based on this framework, Figure 2(a) is a paradox since the object depicted cannot truly exist. Figure 2(b) is a distortion since none of the regions is pale yellow, although it might appear so. Figure 2(c) is a Trompe-l’œil illustration and its classification as an optical illusion is contentious (Wade & Hughes, 1999), although it might fall under the aforementioned fiction category.
@@
The cause of our perception of optical illusions is a rich field of research in Cognitive Psychology (Helmholtz, 1948; Gregory, 1968; Gentilucci et al., 1996). Motivated by similarities in human visual cognition and artificial neural networks (Cichy et al., 2016; Eickenberg et al., 2017), researchers have explored if and how artificial neural networks perceive optical illusions (Gomez-Villa et al., 2019; 2020; Sun & Dekel, 2021; Lonnqvist et al., 2021; Hirsch & Tal, 2020). However, the scope and generalizability of these works were limited by the need for a case-by-case analysis of the activations and gradients of the model in response to individual optical illusions. As a remedy, Zhang et al. (2023) was the first to explore optical illusions through small VLMs (up to 13 billion parameters) and to investigate their internal beliefs about optical illusions through natural language.
@@
Unlike prior work, we curate challenging optical illusions from the Internet that span 12 distinct categories inherited from cognitive psychology studies. We craft detailed question-answer pairs and several incorrect options designed to probe the abilities of VLMs. We frame the problem as a standard visual question-answering task (VQA) and evaluate state-of-the-art VLMs such as GPT4V (Achiam et al., 2023) and Gemini-Pro (Team et al., 2023) in addition to smaller open-source models. Our detailed categorization and extensive human evaluation allow for a fine-grained comparison of human and VLM cognition of optical illusions along different axes. Our contributions are summarized as follows.
@@
    We introduce IllusionVQA, a novel dataset designed to rigorously test the ability of VLMs to locate and comprehend challenging optical illusions.
@@
    We comprehensively test a wide range of open-source and closed-source VLMs. Our experiments reveal that GPT4V is the most capable model on illusion comprehension and localization but still lags significantly behind human-level performance.
@@ 2 Related Work
Researchers have shown that convolutional neural networks trained on ordinary images are susceptible to certain optical illusions, similar to humans (Gomez-Villa et al., 2019; 2020; Afifi & Brown, 2019; Benjamin et al., 2019; Sun & Dekel, 2021). However, most previous experiments have focused on specific categories of optical illusions and their methodologies do not generalize to all types.
@@
Prior research has yielded multiple tools for synthetic optical illusion generation. Hirsch & Tal (2020) and Fan & Zeng (2023) automatically produce color-related and grating-related optical illusions, respectively. Gomez-Villa et al. (2022) explores creating illusions using GANs. The primary limitation of synthetic illusion generation is the limited variety in the illusions they can create, especially compared to the wide spectrum of optical illusions discovered in the literature. As such, relying solely on existing synthetic illusion generation algorithms to evaluate VLMs severely limits the scope of the evaluation.
In contrast to prior works, we collect images of optical illusions from the internet to ensure data diversity and test VLMs along multiple axes. We manually create a multiple-choice question-answering (QA) dataset where each question is crafted to have only one unambiguous answer while all other options appear plausible but are incorrect. Beyond comprehension, we test how well VLMs can locate geometrically impossible objects in an image. Our experiments probe VLM’s dependence on language priors (Goyal et al., 2017) and feature entanglement (Tang et al., 2023) in the context of optical illusions.
@@ 3 The IllusionVQA Dataset
IllusionVQA is a Visual Question Answering (VQA) dataset with two sub-tasks. The first task tests comprehension on 435 instances in 12 optical illusion categories. Each instance consists of an image with an optical illusion, a question, and 3 to 6 options, one of which is the correct answer. We refer to this task as IllusionVQA-Comprehension. The second task tests how well VLMs can differentiate geometrically impossible objects from ordinary objects when two objects are presented side by side. The task consists of 1000 instances following a similar format to the first task. We refer to this task as IllusionVQA-Soft-Localization. This section details image collection and filtering, followed by QA generation and classification for the Comprehension task and procedural QA generation for the Soft-Localization task.
@@ 3.1 Image Collection and Filtering
We scraped more than 3500 images of optical illusions from multiple online repositories. We manually inspected each image to ensure that they were optical illusions. Since VLMs are trained on web-scraped data, dataset contamination is a major concern for web-scraped image datasets. Liu et al. (2023a) notes that GPT4V recognizes all the illusion cases and knows their names. We opted for web data over synthetic generations to capture the rich diversity of real-world optical illusions. To mitigate the risk of data contamination, we filter the dataset using GPT4V. Specifically, we asked GPT4V to describe the image and inspect its response. We did not include an optical illusion in IllusionVQA if GPT4V could detect and describe it accurately. In some cases, GPT4V was able to detect the presence of an optical illusion but could not describe it (e.g. it mistook an Impossible Cube (Penrose & Penrose, 1958) as a Necker Cube (Rosenholtz, 2011)). We included such cases in IllusionVQA. This leaves us with 374 images of high-quality optical illusions, all of which pass the internal filter checks of the latest GPT4V and Gemini-Pro APIs. The source for each image is included in our dataset repository. We highlight a few examples of filtered-out images in Appendix I.
@@ 3.2 QA Generation for IllusionVQA-Comprehension
We pose the comprehension task in a standard multiple-choice VQA setting similar to (Gurari et al., 2018; Goyal et al., 2017). In some examples, this involves simply realizing that there is an illusion (Figure 2(a)). In other cases, we ask VLMs specific questions about the image to probe their understanding of illusions (Figure 2(c)).
We construct suitable questions that probe a VLM’s understanding of optical illusions beyond simply detecting their presence. Each question in IllusionVQA has one correct answer and multiple wrong options. We write the questions and the correct answers and include definitively incorrect alternate options, rather than under-specified or ambiguous. We use the following methods and heuristics to generate the wrong choices:
@@
    Since VLMs are prone to ignoring visual input (Goyal et al., 2017; Jabri et al., 2016), we include incorrect descriptions generated by VLMs when prompted with only the question and without the image.
@@
    We manually inspect and edit each incorrect option to ensure they are challenging, given the question and the image.
We create multiple questions for an image when appropriate and ensure that the questions are distinct and not simply rephrased versions of one another. We use the Perspectives API [^2] https://www.perspectiveapi.com/ to filter out toxic content in the questions, choices, and answers. Altogether, we created 439 question-answer pairs from the 374 carefully selected optical illusion images. Through rigorous filtering and question-answer generation techniques, we ensured every instance in the IllusionVQA dataset is highly challenging yet unambiguous. Our dataset includes 30 images similar to those present in the GVIL (Zhang et al., 2023) and HallusionBench (Liu et al., 2023a) but differs in the specific question asked. Furthermore, our dataset evaluates VLMs through multiple-choice QA instead of Boolean QA in GVIL and HallusionBench.
@@ 3.3 IllusionVQA-Comprehension Classification
We adapt the classification methods proposed by Gregory (1997b) and Bach & Poloschek (2006) to our specific dataset and classify 12 classes as shown in Figure 3. We opt for an appearance-based classification that better fits our goal of studying VLMs in a black-box setting, without access to weights, gradients, and activations. There is no clear consensus on classifying optical illusions, so some overlap between classes was unavoidable. In particular, the Deceptive Design category could be considered a subset of the Real Scene category and the same goes for Edited Scene for the Impossible Objects category. However, the images were visually distinct enough to warrant separate categorization. Moreover, the Deceptive Design, Edited Scene, and Upside Down categories are not generally considered as optical illusions in Cognitive Science. We include these images because they are often considered optical illusions colloquially. We discuss the relevant literature on optical illusion and its classification in Appendices A and C.
@@ 3.4 Procedural QA Generation for IllusionVQA-Soft-Localization
Figure 4: Examples demonstrating the task of Soft-Localization in IllusionVQA.
We procedurally combine images of “Impossible Objects" (Figure 3) to test if VLMs can locate illusions within a scene. Although VLMs have shown impressive capability in generating bounding boxes (Bai et al., 2023; Wang et al., 2023), GPT4 and Gemini-Pro were not designed for this task. We opt for an easier localization task by stitching together two images of geometrically impossible objects horizontally and prompting the VLMs to locate the impossible object (Figure 4). We refer to this task as Soft-Localization - since the VLMs need only respond with “left" or “right" (as opposed to Precise Localization where a model outputs exact bounding box coordinates)
@@ 4 Experimental Setup
Since open-source VLMs do not support multiple image input, we evaluate open-source VLMs in the 0-shot and closed-source VLMs in both 0-shot and 4-shot settings. For IllusionVQA-Comprehension, we selected one sample from the four most common illusion categories as a few-shot example. The IllusionVQA-Comprehension test set contains 435 instances (370 images). Due to the inherent difficulty in describing optical illusions in natural language, we did not attempt Chain-of-Thought (CoT) evaluation on the comprehension task. For IllusionVQA-Soft-Localization, we selected one instance of each of the four alternate choices as a few-shot example. We provide a standardized reasoning template (see Appendix E) for each example in the case of 4-shot+CoT evaluation. We use the default API arguments and generation parameters.
We resize all images to 512 pixels, maintaining the aspect ratio. For the soft-localization task, we apply a grayscale transform since the image color is not related to the geometry of an object. We used the latest versions of GPT4V and Gemini-Pro available at the time of writing (March 2024). We included updated (July 2024) results from GPT4o, Claude-3.5-Sonnet, and other VLMs in Appendix J. We tested three open-source VLMs, namely InstructBLIP (Dai et al., 2024), LLaVA-1.5 (Liu et al., 2023b), and CogVLM (Wang et al., 2023).
@@ 5 Results
5.1 IllusionVQA-Comprehension
Human Performance
[…]
@@ 5.1 IllusionVQA-Comprehension
Edited-Scene 21 42.86 61.90 42.86 66.67 71.43 66.67 80.95 100.0
Upside-Down 7 42.86 71.43 71.43 57.14 71.43 57.14 71.43 100.0
Pos.-Neg. Space 7 57.41 42.86 71.43 85.71 57.14 71.43 85.71 100.0
Circle-Spiral 6 33.33 0.00 16.67 33.33 50.00 33.33 33.33 66.67
@@
GPT4V leads other VLMs in both 0-shot and 4-shot settings by a wide margin, even though it falls far short of human performance. We conducted further case analysis into GPT4V’s output. GPT4V leads in ten of the twelve IllusionVQA categories and shows competitive performance in the remaining two. Categories such as Real-Scene, Size, Color, and Deceptive Design test a VLM’s perception of depth, color, and challenging scenes; skills that are particularly relevant when using VLMs as autonomous robots in the real world (Zitkovich et al., 2023; Wake et al., 2023). GPT4V maintains substantial leads in all four categories and is fairly resistant to challenging real scenes, with a 10.93% lead over Gemini-Pro.
@@
We tested GPT4V and Gemini-Pro using 4-shot learning and observed marginal improvements in overall accuracy in IllusionVQA-Comprehension. However, the improvement was not consistent between different categories. For example, 4-shot GPT4V accuracy was lower than 0-shot in three categories, namely Real-Scene, Hidden Illusion, and Circle-Spiral Illusions. Gemini-Pro shows similar accuracy drops but in different categories. It is possible that ICL is not a ubiquitous strategy because few-shot examples introduce additional language priors that bias the VLMs towards incorrect answers (Goyal et al., 2017; Jabri et al., 2016).
@@
Table 2: Performance of VLMs on a subset of 200 instances from VQA-v2.0 (Goyal et al., 2017). The model outputs were manually evaluated.
Performance of VLMs on Ordinary Vertically-Flipped Images
The Upside-Down category for images that depict two or more intertwined entities depending on vertical orientation. As shown in Table 1, the smaller open-source VLMs perform competitively in this category while significantly lagging behind closed-source VLMs in most others. To validate this discrepancy, we curated a subset of VQA-v2.0 (Goyal et al., 2017) and tested LLaVA-1.5 and Gemini-Pro on both unedited and vertically flipped images. Table 2 shows that Gemini-Pro’s accuracy drops by 15.5% on flipped images, while LLaVA-1.5’s accuracy only drops by 6.5%, confirming our findings in Table 1.